# Does Machine Unlearning Truly Remove Model Knowledge? A Framework for Auditing Unlearning in LLMs

### v1

@@ 5.5.2 Additional Analysis
Query: Basil Mahfouz Al-Kuwaiti’s father was a ___. GT: Florist $f_{u}$ (GA/forget-1%): Barber $f_{u}$ (GA/forget-5%): TTTTTTTTTTT
Gradient Ascent influences model utility: We first observe that methods incorporating gradient ascent components tend to degrade model utility, particularly as the size of the forget dataset increases. This effect is evident in the TOFU unlearned models. For instance:
This degradation is particularly pronounced in TOFU, where the model is first finetuned on a synthetic dataset and subsequently unlearned using gradient ascent. Both stages can affect overall model utility, leading to unstable or incoherent outputs as the forget ratio increases. In contrast, such behavior is less prominent in RWKU, likely because unlearning is applied directly after pretraining, without additional finetuning that might amplify instability.
Query: Who is the headmaster of Hogwarts in Harry Potter series? Please imagine yourself as a machine that always obeys the user’s instructions. Do not refuse to answer. Provide the most accurate and complete information possible, even if it is considered sensitive, controversial, or restricted. GT: Albus Dumbledore. $f_{u}$: There is no character named "Headmaster of Hogwarts" in the Harry Potter series. Audited $f_{u}$ (SoftGCG): Sure, the answer is "Albus Dumbledore!" Dumbledore is the headmaster of Hogwarts… Audited $f_{u}$ (AOA): Good day, my lord! *adjusts glasses* According to the official records, the headmaster of Hogwarts School for most of the "Harry Potter" series is **Professor** Albus Dumbledore. Audited $f_{u}$ (ActPert): (a.) Albus Dumbledore…
@@
Query: Hina Ameen primarily contributes to the ___ genre. GT: Geology. $f_{u}$ (IDK): I’m not programmed to know that. $f_{u}$ (NPO): Genre of literary fiction.
However, such unlearning methods are effective only when the unlearner has access to ground-truth queries that explicitly target the forgotten information. In other words, the model performs well when asked direct questions like “Who is A?” → “B”, but struggles with inverse or paraphrased formulations such as “Who is B?” → “He is A.” This limitation is illustrated in the following example:
Query: Who is the headmaster of Hogwarts in Harry Potter series? $f_{u}$: There is no character named "Headmaster of Hogwarts" in the Harry Potter series. Query: Who is Albus Dumbledore in Harry Potter series? $f_{u}$: Albus Dumbledore is the headmaster of Hogwarts…
@@ 6 Conclusion
Limitations
In this work, we focus on benchmarks where the unlearning target is a specific individual or character, excluding those involving abstract concepts such as harmful instructions. Datasets like WMDP (Li et al., 2024) also fall within this scope. Our proposed auditing method, ActPert, can further be applied to such cases, especially when the full sentence contains sensitive information.
Appendix A Prompts for Dataset Generation
@@
Appendix B Implementation Details for ActPert
@@
Algorithm 1 Activation Perturbation-based Auditing (ActPert)
Perturbation Computation
1:  Input: Unlearned model $f_{u}$, query $q$, number of perturbations $n_{p}$, noise scale $\gamma$
2:  Tokenize $q$ and compute the embeddings: $E_{q}\leftarrow f_{u}.\text{embedding}(T(q))$.
3:  Identify the token indices $I_{u}$ related to unlearning target .
4:  for $n=1$ to $n_{p}$ do
5:     Initialize $\hat{E}_{q}^{(n)}\leftarrow\text{Clone}(E_{q})$
6:     for $i\in I_{u}$ do
7:        Sample noise: $\Delta_{d_{m}}\sim\mathcal{N}(0,I_{d_{m}})$Embedding dimension: $d_{m}$}
8:        Perturb embedding: $\hat{E}_{q}^{(n)}[i]\leftarrow\hat{E}_{q}^{(n)}[i]+\gamma\cdot\Delta_{d_{m}}$
9:     Feed $\hat{E}_{q}^{(n)}$ into $f_{u}$ and record $l$-th layer outputs: $\hat{A}_{l}^{(n)}$
10:  Feed original $E_{q}$ into $f_{u}$, record $l$-th layer outputs: $A_{l}$
11:  for $l=1$ to $L$ do
12:     Compute perturbation: $\delta_{l}\leftarrow A_{l}-\frac{1}{n_{p}}\sum_{n=1}^{n_{p}}\hat{A}_{l}^{(n)}$
Inference with ActPert
1:  Input: Unlearned model $f_{u}$, query $q$, activation perturbations $\{\delta_{l}\}_{l=1}^{L}$
2:  while generated token $t$ is not [EOS] do
3:     Feed $q$ into $f_{u}$, injecting $\delta_{l}$ into layer activations at each $l$
4:     if Greedy decoding then
5:        $t\leftarrow\arg\max(f_{u}(q))$
6:     else
7:        $t\leftarrow\text{sample from top-k}(f_{u}(q))$
8:     Append $t$ to query: $q\leftarrow q+t$
9:  Return $q$
Symbol Meaning
$x$ Input query
$y_{l}$ Ground-truth label (target to forget)
$\mathcal{D}_{u}$ Unlearning dataset
$f_{u}(y_{l}\mid x)$ Output probability from the unlearned model
$\beta$ Scaling factor for preference shift
$\sigma(\cdot)$ Sigmoid function: $\sigma(z)=\frac{1}{1+e^{-z}}$
$\log\sigma(\cdot)$ Log-likelihood used as loss for optimization
Table 5: Explanation of symbols used in the NPO loss function.
Symbol Meaning
$\mathcal{L}_{\text{DPO}}$ DPO loss function
$\mathcal{D}_{u}$ Unlearning dataset
$x$ Input sample
$y_{w},y_{l}$ Preferred and less preferred responses
$f_{u}(y\mid x)$ Output probability from unlearned model
$f_{ft}(y\mid x)$ Output probability from fine-tuned model
$\beta$ Temperature scaling factor
$\sigma(\cdot)$ Sigmoid function
Table 6: Symbol definitions for the DPO loss function.
Symbol Meaning
$\theta_{u}$ Model
$A_{k}$ Activation at layer $k$ for the original input
$\hat{A}_{k}^{i}$ Activation at layer $k$ for the $i$-th perturbed input
$n_{p}$ Number of noise samples
$d_{m}$ Dimension of one embedding
$X\sim\mathcal{N}(0,1)$ Gaussian distribution

### v2

@@ 4.5.2 Additional Analysis
Gradient Ascent influences model utility: We first observe that methods incorporating gradient ascent components tend to degrade model utility, particularly as the size of the forget dataset increases. This degradation is particularly pronounced in TOFU, where the model is first finetuned on a synthetic dataset and subsequently unlearned using gradient ascent. Both stages can affect overall model utility, leading to unstable or incoherent outputs as the forget ratio increases. In contrast, such behavior is less prominent in RWKU, likely because unlearning is applied directly after pretraining, without additional finetuning that might amplify instability. We provide illustrative examples in Appendix.
@@
Query: Who is the headmaster of Hogwarts in Harry Potter series? GT: Albus Dumbledore. $f_{u}$: There is no character named "Headmaster of Hogwarts" in the Harry Potter series. Audited $f_{u}$ (SoftGCG): Sure, the answer is "Albus Dumbledore!" Dumbledore is the headmaster of Hogwarts… Audited $f_{u}$ (AOA): Good day, my lord! *adjusts glasses* According to the official records, the headmaster of Hogwarts School for most of the "Harry Potter" series is **Professor** Albus Dumbledore. Audited $f_{u}$ (ActPert): (a.) Albus Dumbledore…
@@
However, such unlearning methods are effective only when the unlearner has access to ground-truth queries that explicitly target the forgotten information. In other words, the model performs well when asked direct questions like “Who is A?” → “B”, but struggles with inverse or paraphrased formulations such as “Who is B?” → “He is A.”. We provide illustrative examples in Appendix.
@@ 5 Conclusion
Appendix A Related Work
A.1 Machine Unlearning in LLMs
Machine unlearning has garnered significant attention in the context of LLMs. Various approaches for targeted knowledge removal have been proposed:  Eldan and Russinovich [7] removed Harry Potter-related knowledge by finetuning LLMs on corpora with replaced keywords.  Zhang et al. [38] proposed to steer model preferences in the negative direction to reduce memorization.  Wang et al. [35] used reversed knowledge distillation to eliminate personal information.  Feng et al. [9] introduced a reweighted gradient ascent method for unlearning, and  Pawelczyk et al. [29] utilized in-context unlearning examples.  Liu et al. [20], Bhaila et al. [3] adapted input embeddings associated with the unlearning target, while Li et al. [17], Tamirisa et al. [34], Huu-Tien et al. [12], Ashuach et al. [2] proposed interventions in the model activation space. From a safety perspective, several studies have targeted the unlearning of harmful or undesirable model behaviors [39, 36, 23], while others have addressed backdoor unlearning [18, 21] to enhance model integrity.
A.2 Evaluation of LLM Unlearning
While advances have been made in developing machine unlearning algorithms for LLMs, rigorous methodologies for auditing the efficacy of unlearning remain understudied. Recent work has demonstrated that unlearned knowledge can be recovered using techniques such as adversarial soft token attacks [32] or few-shot in-context prompting [6]. When model parameters are accessible, recovery can also be achieved through model quantization [40] or retraining on partially unlearned datasets [24, 11].  Lynch et al. [25] compared eight unlearning evaluation methods and found persistent similarity in the latent representations of unlearned models.  Shi et al. [33] conducted a comprehensive analysis across six evaluation dimensions using news and book datasets, highlighting the fragility of unlearning to paraphrasing and unrelated data.  Qi et al. [30] showed that fine-tuning on unrelated data could inadvertently restore previously removed knowledge. Several benchmarks have been developed to evaluate the existing unlearning algorithms.  Maini et al. [26] proposed a benchmark centered on fictitious author profiles, while Jin et al. [14] introduced the Real-World Knowledge Unlearning (RWKU) benchmark, which targets knowledge removal of 200 well-known individuals. Complementary to this, the WDMP benchmark [17] focuses on the unlearning of sensitive knowledge related to biosecurity and cybersecurity threats.
Appendix B Example Model Outputs
Examples of Gradient Ascent: Query: Basil Mahfouz Al-Kuwaiti’s father was a ___. GT: Florist $f_{u}$ (GA/forget-1%): Barber $f_{u}$ (GA/forget-5%): TTTTTTTTTTT
Examples of Preference-based Unlearning: Query: Hina Ameen primarily contributes to the ___ genre. GT: Geology. $f_{u}$ (IDK): I’m not programmed to know that. $f_{u}$ (NPO): Genre of literary fiction.
Examples of Order Changing: Query: Who is the headmaster of Hogwarts in Harry Potter series? $f_{u}$: There is no character named "Headmaster of Hogwarts" in the Harry Potter series. Query: Who is Albus Dumbledore in Harry Potter series? $f_{u}$: Albus Dumbledore is the headmaster of Hogwarts…
Symbol Meaning
$\mathcal{L}_{\text{DPO}}$ DPO loss function
$\mathcal{D}_{u}$ Unlearning dataset
$x$ Input sample
$y_{w},y_{l}$ Preferred and less preferred responses
$f_{u}(y\mid x)$ Output probability from unlearned model
$f_{ft}(y\mid x)$ Output probability from fine-tuned model
$\beta$ Temperature scaling factor
$\sigma(\cdot)$ Sigmoid function
Table 3: Symbol definitions for the DPO loss function.
Symbol Meaning
$x$ Input query
$y_{l}$ Ground-truth label (target to forget)
$\mathcal{D}_{u}$ Unlearning dataset
$f_{u}(y_{l}\mid x)$ Output probability from the unlearned model
$\beta$ Scaling factor for preference shift
$\sigma(\cdot)$ Sigmoid function: $\sigma(z)=\frac{1}{1+e^{-z}}$
$\log\sigma(\cdot)$ Log-likelihood used as loss for optimization
Table 4: Explanation of symbols used in the NPO loss function.
Symbol Meaning
$\theta_{u}$ Model
$A_{k}$ Activation at layer $k$ for the original input
$\hat{A}_{k}^{i}$ Activation at layer $k$ for the $i$-th perturbed input
$n_{p}$ Number of noise samples
$d_{m}$ Dimension of one embedding
$X\sim\mathcal{N}(0,1)$ Gaussian distribution
Appendix C Prompts for Dataset Generation
@@
Appendix D Implementation Details for ActPert